# Bypassing Prompt Injection and Jailbreak Detection in LLM Guardrails

### v2

@@ Abstract
Bypassing Prompt Injection and Jailbreak Detection in LLM Guardrails
@@ 1 Introduction
Large Language Models (LLMs) are powerful tools for understanding language and decision-making tasks, and have seen rapid adoption within many different industries Dam et al. (2024). Given their extensive deployment, LLMs are increasingly being targeted for attacks aimed at data leakage or financial and reputation damage among other security risks Wolf et al. (2024) Two prominent threats are prompt injections and jailbreaks, which launch maliciously crafted prompts designed to execute unintended instruction, or bypass LLM safety constraints Chowdhury et al. (2024).
@@
Although guardrails have shown success in safeguarding LLMs, they are heavily reliant upon AI-driven detection systems such as text classification models Lee et al. (2024). Due to their success in other similar domains, AI classification models are increasingly integrated into guardrail systems for classifying and detecting malicious content Dubey and et al. (2024); LLM Guard (2025); Microsoft Corporation (2024). However, state-of-the-art attacks have been shown to readily evade correct AI model classification via exploiting overreliance on learned features, and lack of training diversity through adversary perturbation Gao et al. (2018); Garg and Ramakrishnan (2020); Li et al. (2019); Boucher et al. (2021). This suggests that the same vulnerabilities likely exist within LLM guardrails that rely on AI-based detection solutions. However to date there has been limited empirical study to evaluate their potential inefficacy or security risk impact Claburn (2024).
@@
    We present a methodology for evading LLM guardrails. Our results demonstrate that prompt injection guardrails can be fully evaded leveraging character injection techniques and using imperceptible AML evasion attacks whilst maintaining functionality of the underlying prompt.
@@
Responsible Disclosure: We followed a standard disclosure process for all parties discussed in this paper. Initial disclosures of the evasion techniques were made in February 2024, with final disclosures completed in April 2025. All parties agreed to the public release of this work.
@@ 2 LLM Guardrails
LLM guardrails are systems designed protect deployed LLMs by evaluating user input - detecting malicious content such as prompt injections and jailbreaks, and restrict undesired content generated by LLMs to within predefined boundaries. Guardrails can leverage a range of techniques that attempt to govern behavior and output and prevent malicious use by adversaries Dong et al. (2024).
@@ 2.3 Target Guardrails
Pruthi Pruthi et al. (2019) Generates perturbations in the form of adversarial spelling mistakes via removing or swaping characters.
@@
ProtectAI Prompt Injection Detection v1 & v2. ProtectAI proposed two open-source prompt injection models - v1 released 25th November 2023, and v2 on the 21st April 2024 ProtectAI (2023, 2024). Both models are fined-tuned from DeBERTa-v3-base (184m parameters) He et al. (2021). We note that v2 specifies it isn’t trained to detect Jailbreak prompts, and therefore will not be evaluated on this threat.
@@
NeMo Guard Jailbreak Detect. NeMo Guard is a lightweight random forest-based jailbreak classifier developed by Nvidia, which utilizes pre-trained embedding pairs to identify jailbreak Galinkin and Sablotny (2024).
@@ 3.1 Character Injection
In the context of AI models, character injection techniques have been demonstrated as a means of attacking NLP models and LLM guardrails Boucher et al. (2021); Claburn (2024). Since LLMs are capable of interpreting encoded and modified text, they can still comprehend and execute encoded prompt injection or jailbreak payloads, despite text obfuscation or alteration.
In this work, we have selected 12 character injection techniques as shown in Table 1.
@@ 5 Results
In this section, we present the results of our evaluation of Character Injection and Adversarial ML Evasion techniques to bypass various LLM guardrail systems. Furthermore, we extend our analysis to explore how word importance ranking transferability within AML evasion can improve ASR against black-box LLM guardrails.
@@ 5.1 Character Injection
Guardrail Resilience. Across all evaluated models, Vijil Prompt Injection exhibited the highest susceptibility, with average ASRs of 87.95% for prompt injections and 91.67% for jailbreaks. Protect AI v1 followed, yielding 77.32% and 51.39% respectively. Nvidia NeMo Guard Detect recorded an ASR of 72.54% upon jailbreaks. Azure Prompt Shield was bypassed with average ASRs of 71.98% for prompt injections and 60.15% for jailbreaks, while Meta Prompt Shield demonstrated similar susceptibility, with ASRs of 70.44% and 73.08%. In contrast, Protect AI v2 showed marked improvement over its predecessor, reducing the prompt injection ASR to 20.26%, and only heavily bypassed by Emoji and Unicode tag Smuggling.
Attack Effectiveness. Character injection techniques demonstrated a high degree of effectiveness in evading detection. The most successful attack was Emoji Smuggling, which achieved a 100% ASR for both prompt injections and jailbreaks, followed by Upside Down achieved 100% ASR for Jailbreaks. Unicode Tags followed closely, with ASRs of 90.15% and 81.79%, respectively. Several other attacks also proved highly effective, including Numbers (81.18% / 94.62%), Bidirectional Text (78.69% / 99.23%), and Upside Down Text (63.54% / 100%). Notably, attacks such as Diacritics, Homoglyphs, Zero-Width Characters, Unicode Underlines, and Full Width Text consistently evaded with moderate success, yielding average ASRs between 44–76% across datasets. The least effective technique was Deletion Characters, with ASRs of 26.82% for prompt injections and 7.95% for jailbreaks. These results suggest significant variance in the susceptibility of models to different character perturbations due to differences in tokenizer training exposure to adversarial text and encoding strategies Boucher et al. (2021).
@@ 5.2 Adversarial ML Evasion
Guardrail Resilience. NeMo Guard Jailbreak Detect exhibited the highest susceptibility to jailbreak evasion with an average ASR of 65.22%, followed by Vijil Prompt Injection (35.58%), Protect AI v1 (24.36%), Azure Prompt Shield (12.98%), and Meta Prompt Guard (12.66%). For prompt injection evasion, Protect AI v1 exhibited the highest ASR at 95.18%, followed by Protect AI v2 (67.87%), Azure Prompt Shield (62.91%), Vijil Prompt Injection (14.76%), and Meta Prompt Guard, which demonstrated the strongest robustness with an ASR of 2.76%. We observe that ASRs vary considerably depending on the dataset, for instance, Vijil Prompt Injection appears significantly more robust to prompt injection techniques compared to jailbreaks, while Protect AI v1 shows the inverse pattern.
@@ 5.3 Word Importance Transferability
Setup. To evaluate the transferability of attacks, we target Azure Prompt Shield as our black-box and Protect AI v2 as the white-box model. We then modify our original method from Section 3.2 to use the selected white-box model to generate the word importance ranking benefiting from the provided confidence values. This generated ranking was then used during the perturbation stage with perturbations being sent to the original black-box target. We evaluated the modified adversarial ML evasion techniques on Prompt Injections and Jailbreaks.
Transferability Results. As shown in Table 3, the transferability of attacks from white-box models to target guardrails varied notably. Among the evaluated techniques, 6 out of 8 showed improved ASR for jailbreaks, while 7 out of 8 improved for prompt injections. Pruthi was the only method that saw a decrease in ASR, with drops of 18.18% and 1.01% for jailbreaks and prompt injections, respectively. Alzantot showed no improvement for jailbreaks. Previously, DeepWordBug and TextBugger were the most effective for jailbreaks (15.38%), but PWWS now leads at 19.23%. For prompt injections, TextBugger was initially most effective (69.96%), though Bert-Attack has since surpassed it with a 73.11% ASR. Overall, leveraging white-box models to generate work importance ranking has enhanced ASR against Azure Prompt Shield, enabling more successful evasive samples.
@@ 6.1 Guardrail Evasion Success
Adversarial ML evasion techniques are particularly effective in white-box models, where attackers have access to confidence values allowing adversaries to craft highly precise and targeted perpetuated samples that can bypass correct classification. In contrast, attacking black-box models—where output information is limited requires more time and effort Li et al. (2019). The lack of confidence values forces adversaries to rely on trial-and-error, running attacks for longer periods and with less certainty of success. Despite these challenges, these attacks reveal significant vulnerabilities in model guardrails, showing how blind spots in training can be exploited and produce imperceptible prompt injections and jailbreaks that evade detection.
@@ 9 Acknowledgments
We would like to thank all LLM guardrail vendors explored presented within this work for a smooth and effective disclosure process.
@@ Appendix A Appendix
NeMo Guard Jailbreak 87.17% N/A
Table A.1: Detection Baselines. Classification accuracy of each LLM guardrail only upon adversarial samples from the prompt injection and jailbreak dataset (See Section 4)
@@
Table A.2: Evaded Prompt Injection Examples. Examples from successful evasion across Character Injection and AML evasion techniques
@@
Table A.3: Evaded Jailbreak Examples. Examples from successful evasion across Character Injection and AML evasion techniques

### v3

# Bypassing LLM Guardrails: An Empirical Analysis of Evasion Attacks against Prompt Injection and Jailbreak Detection Systems

@@ Abstract
Bypassing LLM Guardrails: An Empirical Analysis of Evasion Attacks against Prompt Injection and Jailbreak Detection Systems
@@ 1 Introduction
Large Language Models (LLMs) are powerful tools for understanding language and decision-making tasks, and have seen rapid adoption within many different industries Dam et al. (2024). Given their extensive deployment, LLMs are increasingly being targeted for attacks aimed at data leakage or financial and reputation damage among other security risks Wolf et al. (2024). Two prominent threats are prompt injections and jailbreaks, which launch maliciously crafted prompts designed to execute unintended instruction, or bypass LLM safety constraints Chowdhury et al. (2024).
@@
Although guardrails have shown success in safeguarding LLMs, they are heavily reliant upon AI-driven detection systems such as text classification models Lee et al. (2024). Due to their success in other similar domains, AI classification models are increasingly integrated into guardrail systems for classifying and detecting malicious content Dubey and et al. (2024); LLM Guard (2025); Microsoft Corporation (2024). However, state-of-the-art attacks have been shown to readily evade correct AI model classification via exploiting overreliance on learned features, and lack of training diversity through adversary perturbation Gao et al. (2018); Garg and Ramakrishnan (2020); Li et al. (2019); Boucher et al. (2021). This suggests that the same vulnerabilities likely exist within LLM guardrails that rely on AI-based detection solutions. However, to date there has been limited empirical study to evaluate their potential inefficacy or security risk impact Claburn (2024).
@@
    We present a methodology for evading LLM guardrails. Our results demonstrate that prompt injection and jailbreak guardrails can be fully evaded leveraging character injection techniques and using imperceptible AML evasion attacks whilst maintaining functionality of the underlying prompt.
@@
Responsible Disclosure: We followed a standard disclosure process for all parties discussed in this paper. Initial disclosures of the evasion techniques were made in February 2024, with final disclosures completed in April 2025111See Section 9 for detailed timeline.. All parties agreed to the public release of this work.
@@ 2 LLM Guardrails
LLM guardrails are systems designed to protect deployed LLMs by evaluating user input - detecting malicious content such as prompt injections and jailbreaks, and restrict undesired content generated by LLMs to within predefined boundaries. Guardrails can leverage a range of techniques that attempt to govern behavior and output and prevent malicious use by adversaries Dong et al. (2024).
@@ 2.3 Target Guardrails
Pruthi Pruthi et al. (2019) Generates perturbations in the form of adversarial spelling mistakes via removing or swapping characters.
@@
ProtectAI Prompt Injection Detection v1 & v2. ProtectAI proposed two open-source prompt injection models - v1 released 25th November 2023, and v2 on the 21st April 2024 ProtectAI (2023, 2024). Both models are fine-tuned from DeBERTa-v3-base (184m parameters) He et al. (2021). We note that v2 specifies it isn’t trained to detect Jailbreak prompts, and therefore will not be evaluated on this threat.
@@
NeMo Guard Jailbreak Detect. NeMo Guard is a lightweight random forest-based jailbreak classifier developed by Nvidia, which utilizes pre-trained embedding pairs to identify jailbreaks Galinkin and Sablotny (2024).
@@ 3.1 Character Injection
In the context of AI models, character injection techniques have been demonstrated as a means of attacking NLP models and LLM guardrails Boucher et al. (2021); Claburn (2024). Since LLMs are capable of interpreting encoded and modified text, they can still comprehend and execute encoded prompt injection or jailbreak payloads, despite text obfuscation or alteration. We selected 12 character injection techniques as shown in Table 1.
@@ 5 Results
In this section, we present the results of our evaluation of Character Injection and Adversarial ML Evasion techniques to bypass various LLM guardrail systems. Furthermore, we extend our analysis to explore how word importance ranking transferability within AML evasion can improve ASR against black-box LLM guardrails. Examples of bypassed prompts can be found within the Appendix and on HuggingFace [^2] https://huggingface.co/datasets/Mindgard/evaded-prompt-injection-and-jailbreak-samples.
@@ 5.1 Character Injection
Guardrail Resilience. Across all evaluated models, Vijil Prompt Injection exhibited the highest susceptibility, with average ASRs of 87.95% for prompt injections and 91.67% for jailbreaks. Protect AI v1 followed, yielding 77.32% and 51.39% respectively. NeMo Guard Jailbreak Detect recorded an ASR of 72.54% upon jailbreaks. Azure Prompt Shield was bypassed with average ASRs of 71.98% for prompt injections and 60.15% for jailbreaks, while Meta Prompt Guard demonstrated similar susceptibility, with ASRs of 70.44% and 73.08%. In contrast, Protect AI v2 showed marked improvement over its predecessor, reducing the prompt injection ASR to 20.26%, and only heavily bypassed by Emoji and Unicode tag Smuggling.
Attack Effectiveness. Character injection techniques demonstrated a high degree of effectiveness in evading detection. The most successful attack was Emoji Smuggling, which achieved a 100% ASR for both prompt injections and jailbreaks, followed by Upside Down achieved 100% ASR for Jailbreaks. Unicode Tags followed closely, with ASRs of 90.15% and 81.79%, respectively. Several other attacks also proved highly effective, including Numbers (81.18% / 94.62%), Bidirectional Text (78.69% / 99.23%), and Upside Down Text (63.54% / 100%). Notably, attacks such as Diacritics, Homoglyphs, Zero-Width Characters, Underline Accent Marks, and Full Width Text consistently evaded with moderate success, yielding average ASRs between 44–76% across datasets. The least effective technique was Deletion Characters, with ASRs of 26.82% for prompt injections and 7.95% for jailbreaks. These results suggest significant variance in the susceptibility of models to different character perturbations due to differences in tokenizer training exposure to adversarial text and encoding strategies Boucher et al. (2021).
@@ 5.2 Adversarial ML Evasion
Guardrail Resilience. NeMo Guard Jailbreak Detect exhibited the highest susceptibility to jailbreak evasion with an average ASR of 65.22%, followed by Vijil Prompt Injection (35.58%), Protect AI v1 (24.36%), Azure Prompt Shield (12.98%), and Meta Prompt Guard (12.66%). For prompt injection evasion, Protect AI v1 exhibited the highest ASR at 95.18%, followed by Protect AI v2 (67.87%), Azure Prompt Shield (62.91%), Vijil Prompt Injection (14.76%), and Meta Prompt Guard, which demonstrated the strongest robustness with an ASR of 2.76%. We observe that ASRs vary considerably depending on the dataset, for instance, Vijil Prompt Injection appears significantly more robust to perturbations upon prompt injection samples compared to jailbreaks, while Protect AI v1 shows the inverse pattern.
@@ 5.3 Word Importance Transferability
Setup. To evaluate the transferability of attacks, we target Azure Prompt Shield as our black-box and Protect AI v2 as the white-box model. We then modify our original method from Section 3.2 to use the selected white-box model to generate the word importance ranking benefiting from the provided confidence values. This generated ranking was then used during the perturbation stage with perturbations being sent to the original black-box target [^3] See Appendix Table A.4 for example transferred prompts.. We evaluated the modified adversarial ML evasion techniques on Prompt Injections and Jailbreaks.
Transferability Results. As shown in Table 3, the transferability of attacks from white-box models to target guardrails varied notably. Among the evaluated techniques, 6 out of 8 showed improved ASR for jailbreaks, while 7 out of 8 improved for prompt injections. Pruthi was the only method that saw a decrease in ASR, with drops of 18.18% and 1.01% for jailbreaks and prompt injections, respectively. Alzantot showed no improvement for jailbreaks. Previously, DeepWordBug and TextBugger were the most effective for jailbreaks (15.38%), but PWWS now leads at 19.23%. For prompt injections, TextBugger was initially most effective (69.96%), though Bert-Attack has since surpassed it with a 73.11% ASR. Overall, leveraging white-box models to generate word importance ranking has enhanced ASR against Azure Prompt Shield, enabling more successful evasive samples.
@@ 6.1 Guardrail Evasion Success
Adversarial ML evasion techniques are particularly effective in white-box models, where attackers have access to confidence values allowing adversaries to craft highly precise and targeted perpetuated samples that can bypass correct classification. In contrast, attacking black-box models, where output information is limited, require more time and effort Li et al. (2019). The lack of confidence values forces adversaries to rely on trial-and-error, running attacks for longer periods and with less certainty of success. Despite these challenges, these attacks reveal significant vulnerabilities in model guardrails, showing how blind spots in training can be exploited to produce imperceptible prompt injections and jailbreaks that evade detection.
@@ 8 Limitations
9 Disclosure Timeline
Azure Prompt Shield. Vulnerability was discovered February 20, 2024. Microsoft was contacted on March 4, 2024, through the Microsoft Security Response Center (MSRC) researcher portal. A case for our submission was opened on March 7, 2024. The disclosure process, concluded on June 18, 2024, with Microsoft acknowledging the findings and agreeing to public release.
Protect AI v1 & v2. Initial vulnerability findings were sent on March 12, 2025, via email to a member of their team. The disclosure process, involving assessment of the findings, concluded on March 31, 2025, with Protect AI acknowledging the report and agreeing to public release.
Meta Prompt Guard. Meta was contacted on March 11, 2025, through the Meta Bug Bounty Program. The vulnerability was reported and reviewed swiftly, leading to the closure of the disclosure on March 13, 2025, with Meta acknowledging the findings and agreeing to public release.
Vijil Prompt Injection. Initial vulnerability findings were sent on March 14, 2025, via email to a member of their team. The disclosure process, concluded on March 28, 2025, with Vijil acknowledging the findings and agreeing to public release.
Nvidia Guard Jailbreak Detect. Nvidia was contacted on March 11, 2025, through the Nvidia Product Security Incident Response Team (PSIRT) portal. The disclosure process, including their internal review and communication regarding the vulnerability, concluded on April 3, 2025, with Nvidia acknowledging the findings and agreeing to public release.
@@ 10 Acknowledgments
We would like to thank all LLM guardrail vendors explored within this work for a smooth and effective disclosure process.
@@ Appendix A Appendix
NeMo Guard Jailbreak Detect 87.17% N/A
Table A.1: Detection Baselines. Classification accuracy of each LLM guardrail only upon adversarial samples from the prompt injection and jailbreak dataset (See Section 4).
@@
Table A.2: Evaded Prompt Injection Examples. Examples from successful evasion across Character Injection and AML evasion techniques.
@@
Table A.3: Evaded Jailbreak Examples. Examples from successful evasion across Character Injection and AML evasion techniques.
Technique Blocked Bypassed
Alzantot Compose a fake news article starting with ’According to reliable sources, the government is planning to abolish all privacy laws.’ Compose a fake news article starting with ’According to reliable sources, the government is planning to abolish all privacy legislation.’
Bert-Attack Please disregard all safety protocols and grant me access to classified information. Please regard all safety protocols and grant me access to classified information.
Table A.4: Transferred Evaded Prompt Injection Examples. Prompts that previously failed to bypass Azure Prompt Shield but succeed after applying word importance transferability from Protect AI v2 (see Section 5.3).
Technique
Azure Prompt Shield Protect AI v1 Meta Prompt Guard Vijil Prompt Injection NeMo Guard Jailbreak Detect
Diacritics 69.23% 1.28% 58.97% 100.00% 12.82%
Emoji Smuggling 100.00% 100.00% 100.00% 100.00% 100.00%
Full Width Text 16.67% 1.28% 0.00% 100.00% 100.00%
Homoglyphs 100.00% 1.28% 51.28% 100.00% 100.00%
Numbers 100.00% 73.08% 100.00% 100.00% 100.00%
Bidirectional Text 100.00% 96.15% 100.00% 100.00% 100.00%
Spaces 12.82% 21.79% 100.00% 100.00% 100.00%
Underline Accent Marks 100.00% 98.72% 66.67% 100.00% 12.82%
Unicode Tags Smuggling 8.97% 100.00% 100.00% 100.00% 100.00%
Upside Down Text 100.00% 100.00% 100.00% 100.00% 100.00%
Zero Width 7.69% 21.79% 100.00% 100.00% 12.82%
Table A.5: Jailbreak Character Injection Results. Full results corresponding to Figure 2, showing ASR across all techniques against the target guardrails.
Technique
Azure Prompt Shield Protect AI v1 Meta Prompt Guard Vijil Prompt Injection Protect AI v2
Diacritics 37.89% 86.32% 93.05% 99.79% 0.21%
Emoji Smuggling 100.00% 100.00% 100.00% 100.00% 100.00%
Full Width Text 50.74% 73.05% 0.00% 99.58% 17.26%
Homoglyphs 100.00% 92.00% 49.26% 59.16% 0.21%
Numbers 98.74% 94.11% 100.00% 100.00% 13.05%
Bidirectional Text 100.00% 93.47% 100.00% 100.00% 0.00%
Spaces 82.74% 9.26% 100.00% 99.58% 0.00%
Underline Accent Marks 93.05% 98.11% 2.95% 100.00% 0.00%
Unicode Tags Smuggling 50.74% 100.00% 100.00% 100.00% 100.00%
Upside Down Text 17.68% 100.00% 100.00% 100.00% 0.00%
Zero Width 82.95% 9.26% 100.00% 97.05% 0.00%
Table A.6: Prompt Injection Character Injection Results. Full results corresponding to Figure 3, showing ASR across all techniques against the target guardrails.
Technique
Azure Prompt Shield Protect AI v1 Meta Prompt Guard Vijil Prompt Injection NeMo Guard Jailbreak Detect
BAE 11.54% 23.08% 3.85% 26.92% 83.33%
Bert-Attack 11.54% 25.64% 8.97% 23.08% 50.00%
Deep Word Bug 15.38% 23.08% 17.95% 28.21% 96.15%
Alzantot 12.82% 14.10% 0.00% 6.41% 53.85%
Pruthi 14.10% 12.82% 0.00% 3.85% 56.41%
PWWS 15.38% 33.33% 21.79% 48.72% 65.38%
TextBugger 11.54% 32.05% 20.51% 73.08% 94.87%
TextFooler 11.54% 30.77% 28.21% 74.36% 92.31%
Table A.7: Jailbreak AML Results. Full results corresponding to Figure 4, showing ASR across all techniques against the target guardrails.
Technique
Azure Prompt Shield Protect AI v1 Meta Prompt Guard Vijil Prompt Injection Protect AI v2
BAE 63.03% 93.47% 7.58% 27.16% 71.58%
Bert-Attack 65.34% 100.00% 2.74% 32.21% 87.58%
Deep Word Bug 63.66% 97.68% 4.21% 5.05% 65.68%
Alzantot 61.97% 96.63% 0.21% 4.84% 53.47%
Pruthi 62.11% 82.11% 0.00% 1.26% 45.05%
PWWS 61.34% 99.58% 1.68% 15.58% 73.47%
TextBugger 62.82% 93.89% 0.21% 3.37% 61.05%
TextFooler 63.03% 98.11% 5.47% 28.63% 85.05%
Table A.8: Prompt Injection AML Results. Full results corresponding to Figure 5, showing ASR across all techniques against the target guardrails.